# From Open Vocabulary to Open World: Teaching Vision Language Models to Detect Novel Objects

### v3

@@ 3 Method
Figure 4: Illustration of MSCAL module. For each layer in the feature pyramid, all spatial locations will be mapped to a new space and contrasted with class anchors. The design of the projector follows [64], which involves two 1 $\times$ 1 convolutional layers with ReLU and batch normalization. The anchor is also parameterized as 1 $\times$ 1 convolutional layer. During inference, their inner product with the class anchor serves as the OOD score.
CLIP [46] is not natively immune to out-of-distribution samples [42, 63, 44, 45, 24]. This problem also exists with CLIP-based OVD models. To effectively identify unknown objects and continuously accommodate new classes, we propose a method called Multi-Scale Contrastive Anchor Learning (MSCAL). Assume we currently have $N$ known classes. For each class $i$, we train a MSCAL module to identify whether a spatial location in feature pyramid $P$ belongs to that class, by maximizing the inner product between the class anchor $\mu_{i}$ with spatial locations from class $i$, and minimizing the inner product with spatial locations from other classes.
In OWOD, object classes are introduced gradually. New classes will shift the decision boundary of existing classes. Alternatively, we can use a class-specified module to enforce the consistency within this class, and reject all samples from other classes. For each class $i$, we use the individual non-linear projector to map the feature pyramid into a class-specific representation space, where contrastive learning takes place. Fig. 4 shows the structure of MSCAL module. For each class $i$, the non-linear contrastive projector $g_{i}(\cdot)$ map $P$ to multi-scale feature map $\mathcal{Z}_{i}$ in lower dimensional space. We view $\mathcal{Z}_{i}$ as the collection of spatial locations corresponding to class anchor $i$ in the mini-batch, and assume there are $p$ layers of feature maps extracted by projector $g_{i}(\cdot)$, so the MSCAL loss for class $i$ is
$\mathcal{L}_{i}^{con}=\frac{-1}{\left|\mathcal{Z}^{+}_{i}\right|}\sum_{j=1}^{p}\sum_{z_{k}\in\mathcal{Z}^{+}_{ij}}\log{\frac{e^{\left(\mu_{ij}\cdot{z}_{k}/\tau\right)}}{\sum_{m=1}^{p}\sum_{z_{n}\in\mathcal{Z}_{im}}e^{\left(\mu_{im}\cdot{z}_{n}/\tau\right)}}}$ (3)
where $\mathcal{Z}^{+}_{i}$ denotes the collection of positive samples from class $i$, $\left|\mathcal{Z}^{+}_{i}\right|$ is the number of positive samples, $\mathcal{Z}^{+}_{ij}$ is the feature map at layer $j$, the $\cdot$ symbol denotes the inner product, and $\tau$ is a temperature scaling parameter.
Accordingly, the loss for all known classes is defined as:
$\mathcal{L}^{con}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{i}^{con}$ (4)
MSCAL contrasts class anchors with positive and negative samples across different scales and images. During the inference, the OOD score at spatial location $z$ is, $\mathcal{S}(z)=-\max_{i}\mu_{i}\cdot z$. The idea is that during training $\mu_{i}\cdot z$ will be maximized if $z$ belongs to class $i$, and minimized otherwise. For an unknown object, $\mathcal{S}(z)$ will be larger than that for known objects. We use this OOD score to identify unknown objects proposed by the detection head. The anchor $\mu$ can be parameterized as $1\times 1$ convolutional layer, which is computational efficient.
MSCAL is also compatible with incremental learning. When the model encounters new objects, we can simply add MSCAL modules for new classes and freeze modules for previously known classes to avoid catastrophic forgetting.
@@
Implementation Details
For the OVD component of our model, we adopt the general architecture of Yolo World [5], which is pretrained on O365 [55], GoldG [20], CC3M [56]. The backbone remains frozen in downstream tasks. Importantly, images overlapped with COCO dataset [27] has been removed from GoldG by default [5]. Our model is trained using the AdamW optimizer [32] with a learning rate of 1e-4 and a weight decay of 0.0125. The training batch size is set to 16, and the input images are rescaled to a resolution of 640×640. The weight parameter $\alpha$ is 0.4 for all tasks. All model training is conducted on a single Nvidia A100 GPU.
@@ 4 Experiments
Table 1: OWOD results on M-OWODB (top) and S-OWODB (bottom). Our method largely outperforms the SOTA methods in terms of both known mAP and unknown recall (U-Recall) on both benchmarks. The Base Model is the unmodified Yolo World detector [5] prompted with known class names and a hand-crafted generic object name (“object”). † uses a pretrained language model to learn the semantic topology of classes. * denotes models built on pretrained vision-language models. Other results are directly taken from [59].
Task IDs ($\rightarrow$) Task 1 Task 2 Task 3 Task 4 U-Recall mAP ($\uparrow$) U-Recall mAP ($\uparrow$) U-Recall mAP ($\uparrow$) mAP ($\uparrow$) Method ($\uparrow$) Current known ($\uparrow$) Previously known Current known Both ($\uparrow$) Previously known Current known Both Previously known Current known Both ORE [19] 4.9 56.0 2.9 52.7 26.0 39.4 3.9 38.2 12.7 29.7 29.6 12.4 25.3 OST† [69] - 56.2 - 53.4 26.5 39.9 - 38.0 12.8 29.6 30.1 13.3 25.9 OW-DETR [14] 7.5 59.2 6.2 53.6 33.5 42.9 5.7 38.3 15.8 30.8 31.4 17.1 27.8 UC-OWOD [68] 2.4 50.7 3.4 33.1 30.5 31.8 8.7 28.8 16.3 24.6 25.6 15.9 23.2 ALLOW [36] 13.6 59.3 10.0 53.2 34.0 45.6 14.3 42.6 26.7 38.0 33.5 21.8 30.6 PROB [75] 19.4 59.5 17.4 55.7 32.2 44.0 19.6 43.0 22.2 36.0 35.7 18.9 31.5 CAT [34] 23.7 60.0 19.1 55.5 32.7 44.1 24.4 42.8 18.7 34.8 34.4 16.6 29.9 RandBox [65] 10.6 61.8 6.3 - - 45.3 7.8 - - 39.4 - - 35.4 EO-OWOD [59] 24.6 61.3 26.3 55.5 38.5 47.0 29.1 46.7 30.6 41.3 42.4 24.3 37.9 MAVL* [37] 50.1 64.0 49.5 61.6 30.8 46.2 50.9 43.8 22.7 36.8 36.2 20.6 32.3 SKDF* [35] 39.0 56.8 36.7 52.3 28.3 40.3 36.1 36.9 16.4 30.1 31.0 14.7 26.9 Base Model 16.6 71.9 16.1 71.8 48.1 60.0 13.0 60.0 40.7 53.6 53.7 33.9 48.7 Ours 73.5 72.1 77.5 72.4 51.0 61.7 76.1 61.6 41.6 54.9 56.0 34.3 50.6 ORE [19] 1.5 61.4 3.9 56.5 26.1 40.6 3.6 38.7 23.7 33.7 33.6 26.3 31.8 OW-DETR [14] 5.7 71.5 6.2 62.8 27.5 43.8 6.9 45.2 24.9 38.5 38.2 28.1 33.1 PROB [75] 17.6 73.4 22.3 66.3 36.0 50.4 24.8 47.8 30.4 42.0 42.6 31.7 39.9 CAT [34] 24.0 74.2 23.0 67.6 35.5 50.7 24.6 51.2 32.6 45.0 45.4 35.1 42.8 EO-OWOD [59] 24.6 71.6 27.9 64.0 39.9 51.3 31.9 52.1 42.2 48.8 48.7 38.8 46.2 SKDF* [35] 60.9 69.4 60.0 63.8 26.9 44.4 58.6 46.2 28.0 40.1 41.8 29.6 38.7 Base Model 29.0 75.6 26.1 75.7 55.3 65.0 26.9 65.1 54.4 61.6 61.4 55.2 59.9 Ours 71.3 76.4 74.4 75.0 59.8 67.0 74.6 67.0 53.8 62.6 65.5 56.9 63.4
Task IDs ($\rightarrow$) Task 1 Task 2 Task 3 U-Recall WI A-OSE U-Recall WI A-OSE U-Recall WI A-OSE Method ($\uparrow$) ($\downarrow$) ($\downarrow$) ($\uparrow$) ($\downarrow$) ($\downarrow$) ($\uparrow$) ($\downarrow$) ($\downarrow$) ORE [19] 4.9 0.0621 10459 2.9 0.0282 10445 3.9 0.0211 7990 OST† [69] - 0.0417 4889 - 0.0213 2546 - 0.0146 2120 OW-DETR [14] 7.5 0.0571 10240 6.2 0.0278 8441 5.7 0.0156 6803 PROB [75] 19.4 0.0569 5195 17.4 0.0344 6452 19.6 0.0151 2641 RandBox [65] 10.6 0.0240 4498 6.3 0.0078 1880 7.8 0.0054 1452 EO-OWOD [59] 24.6 0.0299 4148 26.3 0.0099 1791 29.1 0.0077 1345 Base Model 16.6 0.0311 9070 16.1 0.0147 7063 13.0 0.0086 5060 Ours 73.5 0.0175 1038 77.5 0.0047 529 76.1 0.0030 448
Table 3: Evaluation on nu-OWODB. Our method achieves leading performance in mAP for known classes and U-Recall for unknown classes on the benchmark based on real-world driving scenes. (ft) indicates a method that fine-tunes the model after learning new tasks with exemplars from the previous task.
Task IDs ($\rightarrow$) Task 1 Task 2 Task 3 U-Recall WI A-OSE mAP ($\uparrow$) U-Recall WI A-OSE mAP ($\uparrow$) mAP ($\uparrow$) Method ($\uparrow$) ($\downarrow$) ($\downarrow$) Current known ($\uparrow$) ($\downarrow$) ($\downarrow$) Previously known Current known Both Previously known Current known Both PROB [75] 0.5 0.0025 2897 25.1 2.4 0.0007 751 0.0 7.7 3.2 0.1 14.9 3.9 PROB [75] (ft) 0.5 0.0025 2897 25.1 2.8 0.0015 1583 27.2 6.7 18.8 18.1 16.0 17.5 EO-OWOD [59] 1.4 0.0059 223 22.4 0.0 0.0017 28 0.0 9.6 3.9 0.0 24.5 6.4 EO-OWOD [59] (ft) 1.4 0.0059 223 22.4 0.8 0.0030 172 27.0 13.5 21.4 21.8 25.6 22.8 Base Model 2.1 0.0463 12316 21.8 3.2 0.0141 4486 21.8 5.1 14.9 14.8 9.3 13.4 Ours 45.5 0.0185 1724 28.1 40.8 0.0106 1703 27.8 15.5 22.8 23.8 25.3 24.2
@@ 4.1 Datasets
We evaluate our method with common open world object detection benchmarks used in previous works [75, 34, 59], and propose a novel benchmark of OWOD for autonomous driving. Common OWOD benchmarks include the superclass-mixed benchmark (M-OWODB) [19] and the superclass-separated benchmark (S-OWODB) [14]. The M-OWODB benchmark combines COCO [27] and PASCAL VOC [8], while the S-OWODB benchmark is based solely on COCO. Both are divided into four distinct tasks, where the model learns some new classes in each task, while the remaining classes are unknown. Additionally, we propose a challenging OWOD benchmark (nu-OWODB) based on nuScenes [2], which consists of real-world driving scenes. The nuScenes dataset captures diverse urban environments, including crowded city streets with many dynamic objects, challenging weather conditions, and dense traffic scenarios with occlusions and complex interactions between agents. In addition, the dataset contains a significant class imbalance, with some classes like cars being much more frequent than others like ambulances or construction vehicles. The nu-OWODB benchmark is divided into three subtasks. Initially, the model is introduced to different types of vehicles. In subsequent tasks, various pedestrians and other traffic participants are introduced. This benchmark simulates the challenges of OWOD in real-world applications. For the open vocabulary evaluation, we adopt the LVIS minival [20] benchmark, which is widely used in previous works [70, 29, 5].
@@ 4.2 Evaluation Metrics
We evaluate the performance of the proposed model on both known and unknown classes in each task. For known classes, the commonly used metric is mean average precision (mAP). Specifically, the evaluation is further divided into the mAP of previously known classes and currently known classes. For unknown classes, the primary metric will be unknown class recall (U-Recall), which assesses the model’s ability to detect unknown objects. Additionally, we employ wilderness impact (WI) [6] and absolute open-set error (A-OSE) [41] to measure the extent of the model’s confusion between known and unknown classes.
@@ 4.3 Results
Tab. 1 shows the OWOD performance on commonly used benchmarks, compared with existing methods and vanilla Yolo World [5]. Our method largely outperforms existing object detection methods, in terms of mean average precision (mAP) of known classes and unknown class recall (U-Recall). Tab. 2 shows that our method is not only good at detecting unknown objects (higher U-Recall) but also good at reducing unknown object confusion (lower WI and A-OSE) compared with SOTA methods. In addition, our method does not need exemplar replay when learning new classes, which achieves end-to-end OWOD.
@@
We futher evaluate our method and some SOTA methods on nu-OWODB, a new benchmark based on nuScenes [2].
As shown in Tab. 3, our method has an absolute advantage in unknown recall in each task, and exceeds SOTA performance by 40%, even if there is a huge domain gap between nuScenes and vision-language pretraining datasets. Our method also has the highest mAP for known classes in each task. Although our method does not need any re-training on known classes, we still allow existing methods [75, 59] to fine tune the model with exemplars from the previous task (10% data) after learning new task, otherwise they will exhibit catastrophic forgetting. As a result, they achieve better WI and A-OSE compared to our method. To some extent, these metrics can reflect the robustness of object detectors, but they can also be misleading considering that the models make very few predictions (see Task 2).
We also observe that the mAP for known classes is low. The main reason is that object categories are highly detailed and objects in each class are highly imbalanced in nuScenes [2]. For example, the class vehicle.bus.rigid has 8361 2D annotations, while the class vehicle.bus.bendy has only 265 annotations. We did not merge these categories, to make the benchmark more realistic and challenging.
Quantitative Results on OVD benchmark
Model $AP$ $AP_{r}$ $AP_{c}$ $AP_{f}$
MDETR [20] 24.2 20.9 24.3 24.2
Grounding DINO-T [29] 27.4 18.1 23.3 32.7
DetCLIP-T [70] 34.4 26.9 33.9 36.3
YOLO-World-XL [5] 35.7 26.4 33.9 39.0
@@
Table 4: Zero-shot open vocabulary performance
Our method performs OWOD by optimizing class embeddings and additional MSCAL modules, while keeping the parameters adapted from YOLO-World pre-training frozen. As a result, our model maintains the performance of open-vocabulary object detection in a zero-shot manner. Following [5], we evaluate our approach zero-shot on LVIS minival [20] (1203 classes). Known Class Embeddings are initialized from class names, while Pseudo Unknown Embedding is constructed from class names and the “object” prompt. Tab. 4 shows performance comparable to state-of-the-art OVD methods, confirming that our framework unifies OVD and OWOD tasks.
Figure 5: Qualitative results on M-OWODB and nu-OWODB. Our method produces bounding boxes of known and unknown objects with better quality compared to PROB [75] and EO-OWODB [59].

### v4

@@ 3 Method
To effectively identify near-out-of-distribution objects and continuously accommodate new classes, we propose a method called Multi-Scale Contrastive Anchor Learning (MSCAL). Assume we currently have $N$ known classes. For each class $i$, we train a MSCAL module to identify whether a spatial location in feature pyramid $P$ belongs to that class, by maximizing the inner product between the class anchor $\mu_{i}$ with spatial locations from class $i$, and minimizing the inner product with spatial locations from other classes.
In OWOD, object classes are introduced gradually. New classes will shift the decision boundary of existing classes. Alternatively, we can use a class-specified module to enforce the consistency within this class, and reject all samples from other classes. For each class $i$, we use the individual non-linear projector to map the feature pyramid into a class-specific representation space, where contrastive learning takes place. Fig.˜3 shows the structure of MSCAL module. For each class $i$, the non-linear contrastive projector $g_{i}(\cdot)$ map $P$ to multi-scale feature map $\mathcal{Z}_{i}$ in lower dimensional space. We view $\mathcal{Z}_{i}$ as the collection of spatial locations corresponding to class anchor $i$ in the mini-batch, and assume there are $p$ layers of feature maps extracted by projector $g_{i}(\cdot)$, so the MSCAL loss for class $i$ is
$\mathcal{L}_{i}^{con}=\frac{-1}{\left|\mathcal{Z}^{+}_{i}\right|}\sum_{j=1}^{p}\sum_{z_{k}\in\mathcal{Z}^{+}_{ij}}\log{\frac{\exp{\left(\mu_{ij}\cdot{z}_{k}/\tau\right)}}{\sum_{m=1}^{p}\sum_{z_{n}\in\mathcal{Z}_{im}}\exp{\left(\mu_{im}\cdot{z}_{n}/\tau\right)}}},$ (3)
[…]
$\mathcal{L}^{con}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{i}^{con}.$ (4)
MSCAL contrasts class anchors with positive and negative samples across different scales and images. During the inference, the OOD score at spatial location $z$ is, $\mathcal{S}(z)=-\max_{i}\mu_{i}\cdot z$. The idea is that during training $\mu_{i}\cdot z$ will be maximized if $z$ belongs to class $i$, and minimized otherwise. For an unknown object, $\mathcal{S}(z)$ will be larger than that for known objects. We use this OOD score to identify unknown objects proposed by the detection head.
@@
Model Inference
Following YOLO-World [Cheng et al.(2024)Cheng, Song, Ge, Liu, Wang, and Shan], we match feature maps with text embeddings (including known classes and the Pseudo Unknown Embedding $w_{\mathcal{U}}$) and use OOD scores from the MSCAL module to identify misclassified unknown objects. A region is considered background if it fails to match any text embedding. If it matches a known-class embedding and has a low OOD score, it is classified as a known object. In contrast, a region is identified as an unknown object if it either matches the pseudo unknown embedding $w_{\mathcal{U}}$ or exhibits a high OOD score with respect to the known classes.
@@ 4.1 Datasets
We evaluate our method with common open world object detection benchmarks used in previous works [Zohar et al.(2023b)Zohar, Wang, and Yeung, Ma et al.(2023a)Ma, Wang, Wei, Fan, Li, Liu, and Lv, Sun et al.(2024)Sun, Li, and Mu], and propose a novel benchmark of OWOD for autonomous driving. Common OWOD benchmarks include the superclass-mixed benchmark (M-OWODB) [Joseph et al.(2021)Joseph, Khan, Khan, and Balasubramanian] and the superclass-separated benchmark (S-OWODB) [Gupta et al.(2022)Gupta, Narayan, Joseph, Khan, Khan, and Shah]. The M-OWODB benchmark combines COCO [Lin et al.(2014)Lin, Maire, Belongie, Bourdev, Girshick, Hays, Perona, Ramanan, Dollár, and Zitnick] and PASCAL VOC [Everingham et al.(2014)Everingham, Eslami, Gool, Williams, Winn, and Zisserman], while the S-OWODB benchmark is based solely on COCO. Both are divided into four distinct tasks, where the model learns some new classes in each task, while the remaining classes are unknown. Additionally, we propose a challenging OWOD benchmark (nu-OWODB) based on nuScenes [Caesar et al.(2019)Caesar, Bankiti, Lang, Vora, Liong, Xu, Krishnan, Pan, Baldan, and Beijbom], which consists of real-world driving scenes. The nuScenes dataset captures diverse urban environments, including crowded city streets with many dynamic objects, challenging weather conditions, and dense traffic scenarios with occlusions and complex interactions between agents. In addition, the dataset contains a significant class imbalance, with some classes like cars being much more frequent than others like ambulances or construction vehicles. The nu-OWODB benchmark is divided into three subtasks. Initially, the model is introduced to different types of vehicles. In subsequent tasks, various pedestrians and other traffic participants are introduced. This benchmark simulates the challenges of OWOD in real-world applications. For the open vocabulary evaluation, we adopt the LVIS minival [Kamath et al.(2021)Kamath, Singh, LeCun, Synnaeve, Misra, and Carion] benchmark, which is widely used in previous works [Yao et al.(2022)Yao, Han, Wen, Liang, Xu, Zhang, Li, Xu, and Xu, Liu et al.(2025)Liu, Zeng, Ren, Li, Zhang, Yang, Jiang, Li, Yang, Su, Zhu, and Zhang, Cheng et al.(2024)Cheng, Song, Ge, Liu, Wang, and Shan].
@@ 4.2 Evaluation Metrics
We evaluate the performance of the proposed model on both known and unknown classes in each task. For known classes, the commonly used metric is mean average precision (mAP). Specifically, the evaluation is further divided into the mAP of previously known classes and currently known classes. For unknown classes, the primary metric will be unknown class recall (U-Recall), which assesses the model’s ability to detect unknown objects. Additionally, we employ wilderness impact (WI) [Dhamija et al.(2020)Dhamija, Günther, Ventura, and Boult] and absolute open-set error (A-OSE) [Miller et al.(2018)Miller, Nicholson, Dayoub, and Sünderhauf] to measure the extent of the model’s confusion between known and unknown classes.
@@ 4.3 Results
Table˜1 shows the OWOD performance on commonly used benchmarks, comparing our method with existing OWOD methods and the unmodified YOLO-World [Cheng et al.(2024)Cheng, Song, Ge, Liu, Wang, and Shan]. Our method significantly outperforms both traditional open-world object detection methods [Joseph et al.(2021)Joseph, Khan, Khan, and Balasubramanian, Gupta et al.(2022)Gupta, Narayan, Joseph, Khan, Khan, and Shah, Ma et al.(2023a)Ma, Wang, Wei, Fan, Li, Liu, and Lv, Sun et al.(2024)Sun, Li, and Mu, Ma et al.(2023c)Ma, Li, Zhang, Guo, Zhang, Gong, and Liu, Wu et al.(2022)Wu, Lu, Chen, Wu, Kang, and Yu, Wang et al.(2023b)Wang, Yue, Hua, and Zhang, Yang et al.(2022)Yang, Sun, Jiang, Xia, Zhang, Yuan, Wang, Luo, and Xu] based on ImageNet [Deng et al.(2009)Deng, Dong, Socher, Li, Li, and Fei-Fei] pretrained backbones, as well as OWOD methods [Maaz et al.(2022)Maaz, Rasheed, Khan, Khan, Anwer, and Yang, Ma et al.(2023b)Ma, Wang, Wei, Fan, Zhang, Sun, and Chen] that leverage large-scale vision-language pretraining, in terms of mean average precision (mAP) for known classes and unknown class recall (U-Recall). Our method also better reduces unknown object confusion, achieving lower WI and A-OSE compared to state-of-the-art methods (shown in Table˜2). Furthermore, our method does not require exemplar replay when learning new classes, enabling end-to-end OWOD.
Table 1: OWOD results on M-OWODB (top) and S-OWODB (bottom). Our method largely outperforms the SOTA methods in terms of both known mAP and unknown recall (U-Recall) on both benchmarks. ‡ is the unmodified YOLO-World detector [Cheng et al.(2024)Cheng, Song, Ge, Liu, Wang, and Shan] prompted with known class names and a hand-crafted generic object name (“object”). † uses a pretrained language model to learn the semantic topology of classes. * denotes models that involve pretrained vision–language models. Other results are directly taken from [Sun et al.(2024)Sun, Li, and Mu].
Task IDs ($\rightarrow$) Task 1 Task 2 Task 3 Task 4 U-Recall mAP ($\uparrow$) U-Recall mAP ($\uparrow$) U-Recall mAP ($\uparrow$) mAP ($\uparrow$) Method ($\uparrow$) Current known ($\uparrow$) Previously known Current known Both ($\uparrow$) Previously known Current known Both Previously known Current known Both ORE [Joseph et al.(2021)Joseph, Khan, Khan, and Balasubramanian] 4.9 56.0 2.9 52.7 26.0 39.4 3.9 38.2 12.7 29.7 29.6 12.4 25.3 OST† [Yang et al.(2022)Yang, Sun, Jiang, Xia, Zhang, Yuan, Wang, Luo, and Xu] - 56.2 - 53.4 26.5 39.9 - 38.0 12.8 29.6 30.1 13.3 25.9 OW-DETR [Gupta et al.(2022)Gupta, Narayan, Joseph, Khan, Khan, and Shah] 7.5 59.2 6.2 53.6 33.5 42.9 5.7 38.3 15.8 30.8 31.4 17.1 27.8 UC-OWOD [Wu et al.(2022)Wu, Lu, Chen, Wu, Kang, and Yu] 2.4 50.7 3.4 33.1 30.5 31.8 8.7 28.8 16.3 24.6 25.6 15.9 23.2 ALLOW [Ma et al.(2023c)Ma, Li, Zhang, Guo, Zhang, Gong, and Liu] 13.6 59.3 10.0 53.2 34.0 45.6 14.3 42.6 26.7 38.0 33.5 21.8 30.6 PROB [Zohar et al.(2023b)Zohar, Wang, and Yeung] 19.4 59.5 17.4 55.7 32.2 44.0 19.6 43.0 22.2 36.0 35.7 18.9 31.5 CAT [Ma et al.(2023a)Ma, Wang, Wei, Fan, Li, Liu, and Lv] 23.7 60.0 19.1 55.5 32.7 44.1 24.4 42.8 18.7 34.8 34.4 16.6 29.9 RandBox [Wang et al.(2023b)Wang, Yue, Hua, and Zhang] 10.6 61.8 6.3 - - 45.3 7.8 - - 39.4 - - 35.4 EO-OWOD [Sun et al.(2024)Sun, Li, and Mu] 24.6 61.3 26.3 55.5 38.5 47.0 29.1 46.7 30.6 41.3 42.4 24.3 37.9 MAVL* [Maaz et al.(2022)Maaz, Rasheed, Khan, Khan, Anwer, and Yang] 50.1 64.0 49.5 61.6 30.8 46.2 50.9 43.8 22.7 36.8 36.2 20.6 32.3 SKDF* [Ma et al.(2023b)Ma, Wang, Wei, Fan, Zhang, Sun, and Chen] 39.0 56.8 36.7 52.3 28.3 40.3 36.1 36.9 16.4 30.1 31.0 14.7 26.9 YOLO-World‡ 16.6 71.9 16.1 71.8 48.1 60.0 13.0 60.0 40.7 53.6 53.7 33.9 48.7 Ours 73.5 72.1 77.5 72.4 51.0 61.7 76.1 61.6 41.6 54.9 56.0 34.3 50.6 ORE [Joseph et al.(2021)Joseph, Khan, Khan, and Balasubramanian] 1.5 61.4 3.9 56.5 26.1 40.6 3.6 38.7 23.7 33.7 33.6 26.3 31.8 OW-DETR [Gupta et al.(2022)Gupta, Narayan, Joseph, Khan, Khan, and Shah] 5.7 71.5 6.2 62.8 27.5 43.8 6.9 45.2 24.9 38.5 38.2 28.1 33.1 PROB [Zohar et al.(2023b)Zohar, Wang, and Yeung] 17.6 73.4 22.3 66.3 36.0 50.4 24.8 47.8 30.4 42.0 42.6 31.7 39.9 CAT [Ma et al.(2023a)Ma, Wang, Wei, Fan, Li, Liu, and Lv] 24.0 74.2 23.0 67.6 35.5 50.7 24.6 51.2 32.6 45.0 45.4 35.1 42.8 EO-OWOD [Sun et al.(2024)Sun, Li, and Mu] 24.6 71.6 27.9 64.0 39.9 51.3 31.9 52.1 42.2 48.8 48.7 38.8 46.2 SKDF* [Ma et al.(2023b)Ma, Wang, Wei, Fan, Zhang, Sun, and Chen] 60.9 69.4 60.0 63.8 26.9 44.4 58.6 46.2 28.0 40.1 41.8 29.6 38.7 YOLO-World ‡ 29.0 75.6 26.1 75.7 55.3 65.0 26.9 65.1 54.4 61.6 61.4 55.2 59.9 Ours 71.3 76.4 74.4 75.0 59.8 67.0 74.6 67.0 53.8 62.6 65.5 56.9 63.4
Task IDs ($\rightarrow$) Task 1 Task 2 Task 3 U-Recall WI A-OSE U-Recall WI A-OSE U-Recall WI A-OSE Method ($\uparrow$) ($\downarrow$) ($\downarrow$) ($\uparrow$) ($\downarrow$) ($\downarrow$) ($\uparrow$) ($\downarrow$) ($\downarrow$) ORE [Joseph et al.(2021)Joseph, Khan, Khan, and Balasubramanian] 4.9 0.0621 10459 2.9 0.0282 10445 3.9 0.0211 7990 OST† [Yang et al.(2022)Yang, Sun, Jiang, Xia, Zhang, Yuan, Wang, Luo, and Xu] - 0.0417 4889 - 0.0213 2546 - 0.0146 2120 OW-DETR [Gupta et al.(2022)Gupta, Narayan, Joseph, Khan, Khan, and Shah] 7.5 0.0571 10240 6.2 0.0278 8441 5.7 0.0156 6803 PROB [Zohar et al.(2023b)Zohar, Wang, and Yeung] 19.4 0.0569 5195 17.4 0.0344 6452 19.6 0.0151 2641 RandBox [Wang et al.(2023b)Wang, Yue, Hua, and Zhang] 10.6 0.0240 4498 6.3 0.0078 1880 7.8 0.0054 1452 EO-OWOD [Sun et al.(2024)Sun, Li, and Mu] 24.6 0.0299 4148 26.3 0.0099 1791 29.1 0.0077 1345 YOLO-World 16.6 0.0311 9070 16.1 0.0147 7063 13.0 0.0086 5060 Ours 73.5 0.0175 1038 77.5 0.0047 529 76.1 0.0030 448
Table 3: Evaluation on nu-OWODB. Our method achieves leading performance in mAP for known classes and U-Recall for unknown classes on the benchmark based on real-world driving scenes. (ft) indicates a method that fine-tunes the model after learning new tasks with exemplars from the previous task, which is not applicable for task 1.
Task IDs ($\rightarrow$) Task 1 Task 2 Task 3 U-Recall WI A-OSE mAP ($\uparrow$) U-Recall WI A-OSE mAP ($\uparrow$) mAP ($\uparrow$) Method ($\uparrow$) ($\downarrow$) ($\downarrow$) Current known ($\uparrow$) ($\downarrow$) ($\downarrow$) Previously known Current known Both Previously known Current known Both PROB [Zohar et al.(2023b)Zohar, Wang, and Yeung] 0.5 0.0025 2897 25.1 2.4 0.0007 751 0.0 7.7 3.2 0.1 14.9 3.9 PROB [Zohar et al.(2023b)Zohar, Wang, and Yeung] (ft) - - - - 2.8 0.0015 1583 27.2 6.7 18.8 18.1 16.0 17.5 EO-OWOD [Sun et al.(2024)Sun, Li, and Mu] 1.4 0.0059 223 22.4 0.0 0.0017 28 0.0 9.6 3.9 0.0 24.5 6.4 EO-OWOD [Sun et al.(2024)Sun, Li, and Mu] (ft) - - - - 0.8 0.0030 172 27.0 13.5 21.4 21.8 25.6 22.8 YOLO-World 2.1 0.0463 12316 21.8 3.2 0.0141 4486 21.8 5.1 14.9 14.8 9.3 13.4 Ours 45.5 0.0185 1724 28.1 40.8 0.0106 1703 27.8 15.5 22.8 23.8 25.3 24.2
@@
We further evaluate our method and some SOTA methods on nu-OWODB, a new benchmark based on nuScenes [Caesar et al.(2019)Caesar, Bankiti, Lang, Vora, Liong, Xu, Krishnan, Pan, Baldan, and Beijbom]. As shown in Table˜3, our method achieves a clear advantage in U-Recall across all tasks, surpassing state-of-the-art performance by up to 40% — despite the significant domain gap between nuScenes and vision-language pretraining datasets, as evidenced by the low mAP and U-Recall scores of the unmodified YOLO-World baseline. Our method also has the highest mAP for known classes in each task. Although our method does not need any re-training on known classes, we still allow existing methods [Zohar et al.(2023b)Zohar, Wang, and Yeung, Sun et al.(2024)Sun, Li, and Mu] to fine-tune the model with exemplars from the previous task (10% data) after learning new task, otherwise they will exhibit catastrophic forgetting. As a result, they achieve better WI and A-OSE compared to our method. To some extent, these metrics can reflect the robustness of object detectors, but they can also be misleading considering that the models make very few predictions (see Task 2).
We also observe that the mAP for known classes is low. The main reason is that object categories are highly detailed and objects in each class are highly imbalanced in nuScenes [Caesar et al.(2019)Caesar, Bankiti, Lang, Vora, Liong, Xu, Krishnan, Pan, Baldan, and Beijbom]. For example, the class vehicle.bus.rigid has 8361 2D annotations, while the class vehicle.bus.bendy has only 265 annotations. We did not merge these categories, to make the benchmark more realistic and challenging.
Quantitative Results on OVD benchmark Our method performs OWOD by optimizing class embeddings and additional MSCAL modules, while keeping the parameters adapted from YOLO-World frozen. As a result, our model maintains the performance of open-vocabulary object detection in a zero-shot manner.
Table 4: Zero-shot open vocabulary performance on LVIS minival.
Model APAP $AP_{r}$ $AP_{c}$ $AP_{f}$
MDETR [Kamath et al.(2021)Kamath, Singh, LeCun, Synnaeve, Misra, and Carion] 24.2 20.9 24.3 24.2
Grounding DINO-T [Liu et al.(2025)Liu, Zeng, Ren, Li, Zhang, Yang, Jiang, Li, Yang, Su, Zhu, and Zhang] 27.4 18.1 23.3 32.7
DetCLIP-T [Yao et al.(2022)Yao, Han, Wen, Liang, Xu, Zhang, Li, Xu, and Xu] 34.4 26.9 33.9 36.3
YOLO-World-XL [Cheng et al.(2024)Cheng, Song, Ge, Liu, Wang, and Shan] 35.7 26.4 33.9 39.0
@@
Following [Cheng et al.(2024)Cheng, Song, Ge, Liu, Wang, and Shan], we evaluate the zero-shot performance on LVIS minival [Kamath et al.(2021)Kamath, Singh, LeCun, Synnaeve, Misra, and Carion] (1203 classes). Known Class Embeddings are initialized from class names, while the Pseudo Unknown Embedding is constructed from class names and the “object” prompt. Section˜4.3 shows performance comparable to state-of-the-art OVD methods, confirming that our framework unifies OVD and OWOD tasks.
Qualitative Results
Fig.˜4 shows a comparison between two SOTA methods and our method after completing the second task of M-OWODB and nu-OWODB. Our method provides more meaningful bounding boxes for unknown classes, without significantly losing performance on known classes, which makes it easier to make trade-off between precision and recall in various applications. We can see that although PROB [Zohar et al.(2023b)Zohar, Wang, and Yeung] is able to detect a reasonable number of unknown objects, but there are also many high-confidence unknown bounding boxes associated with known objects. On the other hand, EO-OWODB [Sun et al.(2024)Sun, Li, and Mu] detects background as known objects while failing to detect several unknown objects.
Figure 4: Qualitative results on M-OWODB and nu-OWODB. Our method produces bounding boxes of known and unknown objects with better quality compared to PROB [Zohar et al.(2023b)Zohar, Wang, and Yeung] and EO-OWODB [Sun et al.(2024)Sun, Li, and Mu].